# Assignment–Routing Optimization with Cutting-Plane Subtour Elimination: Solver and Benchmark Dataset

### v1

@@ Abstract
We study a joint routing–assignment optimization problem in which a set of items must be paired one-to-one with a set of placeholders while simultaneously determining a Hamiltonian cycle that visits every node exactly once. Both the assignment and routing decisions are optimized jointly to minimize the total travel cost. In this work, we propose a method to solve this problem using an exact MIP formulation with Gurobi, including cutting-plane subtour elimination. With analysis of the computational complexity and through extensive experiments, we analyze the computational limitations of this approach as the problem size grows and reveal the challenges associated with the need for more efficient algorithms for larger instances. The dataset, formulations, and experimental results provided here can serve as benchmarks for future studies in this research area. GitHub repository : https://github.com/QL-YUAN/Joint-Assignment-Routing-Optimization
@@ 1.1 Connections to existing problem classes
Pickup-and-delivery variants of the traveling salesman and vehicle routing problems (PDTSP, VRP-PD) enforce precedence constraints between pickup and delivery pairs under vehicle capacity limits. The stacker-crane abstraction (Hernández-Pérez and Salazar-González, 2004, 2007; Treleaven et al., 2012) is closely related, representing a single-vehicle system that must transport items between paired pickup and drop-off locations. These problems are NP-hard in general and have motivated a rich literature on branch-and-cut formulations, exact algorithms for small-scale instances, and heuristic or metaheuristic methods for larger ones.
@@
Classical formulations of the assignment problem (Kuhn, 1955) and the traveling salesman problem (Dantzig et al., 1954; Lawler et al., 1985; Applegate et al., 2006) address individual subcomponents of our task. However, when assignment and routing are coupled the search space expands combinatorially, requiring integrated formulations or hierarchical decomposition. Recent robotics literature has explored mixed-integer and sampling-based approaches for such coupled problems, yet scalability and global optimality remain open challenges. Modern solvers such as Gurobi (Gurobi Optimization, LLC, 2024) allow constraint families to be activated dynamically through cutting-plane method based callbacks, a strategy that has been successfully employed in subtour elimination for TSPs and connectivity-enforced routing (Applegate et al., 2006). This mechanism reduces model size and computational overhead by adding connectivity constraints only when violated, which we exploit in our implementation.
@@ 2 Mathematical Formulation
[…]
@@
Constraints are as follows:
@@
    Valid edges: Only item-to-placeholder edges are allowed, and edges can only exist if the assignment exists:
@@
    Subtour elimination: Disconnected cycles are prohibited by a cutting-plane method Kelley (1960); Westerlund and Pettersson (1995):
@@ 3 Computational Complexity
The pure assignment problem can be solved in polynomial time using the Hungarian algorithm Kuhn (1955) in $O(n^{3})$ complexity. However, the present formulation couples the assignment decision with routing optimization, the Traveling Salesman Problem (TSP) with fixed pick-place tasks constraints with complexity of $O(n!)$, making the problem significantly harder in $O(n!*n^{3})$ complexity.
@@ 3.1 Computational Complexity of The Proposed MIP Solver
  • $O(n^{2})$ binary routing variables $x_{ij}$,
  • $O(n^{2})$ binary assignment variables $a_{ip}$,
  • $O(n^{2})$ linear constraints.
  • An exponential number of subtour elimination constraints introduced dynamically by the cutting-plane procedure Basu et al. (2020).
Let $m$ denote the number of binary decision variables in the MIP. For our specific formulation, the dominant binary variables are the routing and assignment binaries:
$m=\Theta(n^{2}),$
At each search instance the solver typically (i) solves a linear programming (LP) relaxation and (ii) runs presolve / separation routines. Let $T_{\text{LP}}(s)$ denote the time to solve an LP of size $s$ (which is $O(2^{m})$) and $T_{\text{sep}}(s)$ denotes the time spent on all separation/cut generation at the instance. LP algorithms are proved to be polynomial in $s$ Karmarkar (1984), so we can write $T_{\text{LP}}(s)=\operatorname{poly}(s)$. However $T_{\text{sep}(s)}$ is exponential function of $s$ in worse-case, according to study in Basu et al. (2020). In other words, in worse-case, the MIP algorithm has exponential worst-case time complexity even for only one single instance. Thus, the worst-case computational effort grows exponentially with problem size ($n$), and only small to medium instances can be solved to proven optimality using current MIP solvers Basu et al. (2020); Optimization (2025).
@@ 4 Datasets
All the datasets files can be downloaded from project GitHub repository.
@@
  • assignment: Selected assignment solution (represented as a list of items assigned to placeholders).
@@
n=32_icn=32\_ic gurobi_n_32_ic_results.csv img_32_ic
n=32_xqn=32\_xq gurobi_n_32_xq_results.csv img_32_xq
$n=100$ gurobi_n_100_results.csv img_100
$n=200$ gurobi_n_200_results.csv img_200
$n=300$ gurobi_n_300_results_1_30.csv img_300
Table 4: Summary of Gurobi results for sample experiments (timeout = 5 min)
@@ 5 Experimental Results
Samples of the results are shown in Figures 1234, and summarized in Table 5. The complete dataset of experimental results is available in the accompanying GitHub repository for full reproducibility.
As evidenced in Table 5, the exact solver performs efficiently for small instances, consistently returning optimal solutions in under a second for problem sizes up to $n=32$. However, as the problem size increases — particularly for $n=300$ (30 sample instances are included in the dataset) — the computational time becomes significantly larger. For most instances at this scale, the solver requires more than 10 minutes to complete, even when parallelized.
@@
Table 5: Merged experimental results for $n=32$ XQ, $n=32$ Chess, $n=100$, and $n=200$ placeholders.
$n=32$ XQ $n=32$ Chess $n=100$ $n=200$
Item Best Cost Time (s) Item Best Cost Time (s) Exp ID Best Cost Time (s) Item Best Cost Time (s)
1 8.9811 0.3522 1 8.6045 0.2179 1 19.5360 9.6141 1 32.5257 25.6686
2 11.1315 0.0791 2 11.4452 0.0763 2 30.7337 10.9660 2 29.7790 206.0344
3 12.5431 0.2851 3 13.5874 0.2403 3 26.1106 3.7165 3 26.9493 47.0434
4 10.6235 0.0617 4 11.2942 0.2280 4 19.0954 24.1544 4 27.4590 49.4273
5 10.9402 0.2697 5 10.8882 0.3152 5 18.5475 7.0274 5 28.5458 124.0022
6 8.5998 0.3594 6 9.5426 2.2543 6 19.2728 14.9589 6 32.5525 78.4429
7 10.8217 0.1051 7 11.3253 0.2189 7 20.7782 20.2498 7 38.1046 95.0898
8 9.3054 0.6133 8 9.8737 0.1141 8 25.1340 2.5392 8 38.0598 45.8259
9 10.6951 0.3497 9 11.3153 0.4692 9 19.9261 6.4724 9 31.5440 65.1356
10 11.4083 0.0931 10 12.3702 0.1388 10 20.6597 10.5358 10 28.0293 106.9303
Average 0.3062 Average 0.3579 Average 11.5 Average 107.7101
@@ 6 Conclusion
In this work, we addressed a joint assignment–routing optimization problem, where item-to-placeholder pairings and a Hamiltonian tour are determined simultaneously to minimize total travel cost. We formulated the problem as a compact mixed-integer program (MIP) and solved it using Gurobi, employing a cutting-plane approach for subtour elimination to ensure tour feasibility.
Our experimental analysis confirms that the exact MIP solver performs well for moderate-sized instances (e.g., $n=32$), typically producing optimal solutions in fractions of a second. However, as the problem size increases, computational requirements grow rapidly: for $n=200$, solving times exceed 100 seconds; for $n=300$, runtimes often exceed 10 minutes per instance. These results highlight the scalability limitations of exact MIP-based methods in practice.

### v2

@@ Abstract
We study a joint routing–assignment optimization problem in which a set of items must be paired one-to-one with a set of placeholders while simultaneously determining a Hamiltonian cycle that visits every node exactly once. Both the assignment and routing decisions are optimized jointly to minimize the total travel cost. In this work, we propose a method to solve this problem using an exact MIP formulation with Gurobi, including cutting-plane subtour elimination. With analysis of the computational complexity and through extensive experiments, we analyze the computational limitations of this approach as the problem size grows and reveal the challenges associated with the need for more efficient algorithms for larger instances. The dataset, formulations, and experimental results provided here can serve as benchmarks for future studies in this research area. Many pickup-placing with routing problems can be solved with such method with addition assignment constraints. GitHub repository : https://github.com/QL-YUAN/Joint-Assignment-Routing-Optimization
@@ 1.1 Connections to existing problem classes
Pickup-and-delivery variants of the traveling salesman and vehicle routing problems (PDTSP, VRP-PD) enforce precedence constraints between pickup and delivery pairs under vehicle capacity limits. The stacker-crane abstraction hernandez2004 ; hernandez2007 ; treleaven2012 is closely related, representing a single-vehicle system that must transport items between paired pickup and drop-off locations. These problems are NP-hard in general and have motivated a rich literature on branch-and-cut formulations, exact algorithms for small-scale instances, and heuristic or metaheuristic methods for larger ones.
@@
Classical formulations of the assignment problem kuhn1955hungarian and the traveling salesman problem dantzig1954solution ; lawler1985traveling ; applegate2006traveling address individual subcomponents of our task. However, when assignment and routing are coupled the search space expands combinatorially, requiring integrated formulations or hierarchical decomposition. Recent robotics literature has explored mixed-integer and sampling-based approaches for such coupled problems, yet scalability and global optimality remain open challenges. Modern solvers such as Gurobi gurobi2024manual allow constraint families to be activated dynamically through cutting-plane method based callbacks, a strategy that has been successfully employed in subtour elimination for TSPs and connectivity-enforced routing applegate2006traveling . This mechanism reduces model size and computational overhead by adding connectivity constraints only when violated, which we exploit in our implementation.
@@ 2 Mathematical Formulation
$x_{ij}=\begin{cases}1,&\text{if edge }(i,j)\text{ is included in the cycle},\\[4.0pt] 0,&\text{otherwise.}\end{cases}$
@@
2.1 Simplified Model Constraints
In order to solve the problem efficiently, in this problem modeling, only edges from item to placeholder is allowed. Constraints are as follows:
  1. 1.
    Valid edges: In order to make the model as simple as possible for high high computational efficiency , here only item-to-placeholder edges are allowed. Since straight forward post processing of edges can result in complete cycle with directions, such definition is fine for standard assignment and routing problem. The valid edges constraints are as follows:
[…]
  2. 2.
    Assignment: Each item and placeholder is assigned twice:
    $\sum_{p\in P}x_{ip}=2,\quad\forall i\in I,\quad\sum_{i\in I}x_{ip}=2,\quad\forall p\in P$
  3. 3.
    $x_{n-1,2n-1}=1,x_{2n-1,n-1}=0$
  4. 4.
    Subtour elimination: Disconnected cycles are prohibited by a cutting-plane method kelley1960cutting ; westerlund1995extended :
2.2 Generalization of Modeling for Problem with More Constraints
The formulation introduced above is suitable for the case of a single type of items and placeholders, where each item can be assigned to any placeholder. However, many practical applications involve multiple types or categories of items, each associated with a distinct subset of compatible placeholders. Examples include multi-type packing tasks, board or game piece arrangements, and robotic placement operations where items of the same type are interchangeable but must be placed in designated regions.
We will have follow-up articals to discuss on methods to solve such problems. However, first of all, in such scenarios, the simplified model as mentioned above will have difficulty defining the constraints. In such cases we would like the edges to be directional instead of only connecting from item to placeholder. At the same time, we need to introduce additional binary decision variables:
$a_{ip}=\begin{cases}1&\text{if item }i\text{ is assigned to placeholder }p\\ 0&\text{otherwise}\end{cases}.$
Then, the constraints of the joint assignment-routing problem becomes:
@@ 2.2 Generalization of Modeling for Problem with More Constraints
    Valid edges: Only item-to-placeholder edges are allowed, and edges can only exist if the assignment exists (here placeholder to item edges is allowed in this modeling):
@@
    Subtour elimination: Disconnected cycles are prohibited by a cutting-plane method:
@@
  6. 6.
    Following constraints need to be added to speed up the solver.
    $x_{ij}+x_{ji}<=1,\forall i,\forall j,i\neq j$
@@ 3 Computational Complexity
The pure assignment problem can be solved in polynomial time using the Hungarian algorithm kuhn1955hungarian in $O(n^{3})$ complexity. However, the present formulation couples the assignment decision with routing optimization, the Traveling Salesman Problem (TSP) with fixed pick-place tasks constraints with complexity of $O(n!)$, making the problem significantly harder in $O(n!*n^{3})$ complexity.
@@ 3.1 Computational Complexity of The Proposed MIP Solver
  • $O(n^{2})$ binary routing variables $x_{ij}$, and $O(n^{2})$ linear constraints.
  • An exponential number of subtour elimination constraints introduced dynamically by the cutting-plane procedure basu2020complexitybranchandboundcuttingplanes .
Let $m$ denote the number of binary decision variables in the MIP. For our specific formulation, the dominant binary variables are the routing and assignment binaries: $m=\Theta(n^{2}),$
At each search instance the solver typically (i) solves a linear programming (LP) relaxation and (ii) runs presolve / separation routines. Let $T_{\text{LP}}(s)$ denote the time to solve an LP of size $s$ (which is $O(2^{m})$) and $T_{\text{sep}}(s)$ denotes the time spent on all separation/cut generation at the instance. LP algorithms are proved to be polynomial in $s$ karmarkar1984new , so we can write $T_{\text{LP}}(s)=\operatorname{poly}(s)$. However $T_{\text{sep}(s)}$ is exponential function of $s$ in worse-case, according to study in basu2020complexitybranchandboundcuttingplanes . In other words, in worse-case, the MIP algorithm has exponential worst-case time complexity. Thus, the worst-case computational effort grows exponentially with problem size ($n$), and only small to medium instances can be solved to proven optimality using current MIP solvers basu2020complexitybranchandboundcuttingplanes ; gurobi2025mipprimer .
@@ 4 Datasets
All the datasets files can be downloaded from project GitHub repository. The experiments presented here applied the simplified model as introduced in Section 2.
@@
n=32_icn=32\_ic updated_experiment_n_32_ic_results.csv img_update_32_ic
n=32_xqn=32\_xq updated_experiment_n_32_xq_results.csv img_update_32_xq
$n=100$ updated_experiment_n_100_results.csv img_update_100
$n=200$ updated_experiment_n_200_results.csv img_update_200
$n=300$ updated_experiment_n_300_results_1_30.csv img_update_updated_300
Table 4: Summary of Gurobi results for sample experiments (n=100)
@@ 5 Experimental Results
Samples of the results are shown in Figures 1234, and summarized in Table 5. The complete dataset of experimental results is available in the accompanying GitHub repository for full reproducibility. As evidenced in Table 5, the exact solver performs efficiently for small instances, consistently returning optimal solutions in under a second for problem sizes up to $n=100$. However, as the problem size increases — particularly for $n=300$, the computational time reach one minutes. When n reach 1000, the solver often fails to provide a valid solution.
@@
Table 5: Merged experimental results for $n=32$ XQ, $n=32$ Chess, $n=100$, $n=200$, and $n=300$ placeholders.
$n=32$ XQ $n=32$ Chess $n=100$ $n=200$ $n=300$
Item Best Cost Time (s) Best Cost Time (s) Best Cost Time (s) Best Cost Time (s) Best Cost Time (s)
1 8.9811 0.1749 8.6045 0.0640 19.5360 0.5980 32.5257 1.8281 42.0017 76.0503
2 11.1315 0.0184 11.4452 0.0204 30.7337 0.7149 29.7790 8.3149 33.1289 25.6637
3 12.5431 0.2018 13.5874 0.0154 26.1106 0.8715 26.9493 4.2116 34.3303 44.0879
4 10.6235 0.0195 11.2942 0.0401 19.0954 0.9859 27.4590 6.2985 35.1968 28.5358
5 10.9402 0.0450 10.8882 0.0494 18.5475 2.0169 28.5458 14.5061 33.8118 77.8469
6 8.5998 0.0611 9.5426 0.0869 19.2728 1.5554 32.5525 6.9940 37.4785 84.7612
7 10.8217 0.0149 11.3253 0.0717 20.7782 0.8501 38.1054 8.7085 33.9530 26.5168
8 9.3054 0.1165 9.8737 0.0133 25.1340 0.2294 38.0598 4.1589 40.6682 28.3229
9 10.6951 0.0742 11.3153 0.0821 19.9261 0.2498 31.5435 10.0047 36.6766 51.3296
10 11.4083 0.0331 12.3702 0.0123 20.6597 2.6301 28.0285 11.0655 38.6581 83.7323
Average 0.0758 0.0452 0.9709 7.6993 52.2849
@@ 6 Conclusion
In this work, we addressed a joint assignment–routing optimization problem, where item-to-placeholder pairings and a Hamiltonian tour are determined simultaneously to minimize total travel cost. We formulated the problem as a compact mixed-integer program (MIP) and solved it using Gurobi, employing a cutting-plane approach for subtour elimination to ensure tour feasibility. Many pickup-placing with routing problems, for example, arranging chess game (with different type of pieces) and packing food/ toys with different item patterns, can be solved with such method with addition assignment constraints without increasing the complexity of the problem.
Our experimental analysis confirms that the exact MIP solver performs well for moderate-sized instances (e.g., $n=30$), typically producing optimal solutions in fractions of a second. However, as the problem size increases, computational requirements grow rapidly: for $n=300$, solving times reach 1 minute; for $n=1000$, solver often fails to work. These results highlight the scalability limitations of exact MIP-based methods in practice.